OpenStudy (anonymous):

What is the probability of rolling a number cube twice and getting an odd number each time?

OpenStudy (zarkon):

1/4

OpenStudy (anonymous):

1/2 the rolls are independent

OpenStudy (zarkon):

\[\frac{1}{2}\times\frac{1}{2}\]

OpenStudy (zarkon):

\[P(\text{odd on }1^{st} \text{ and }\text{odd on }2^{nd})\] \[=P(\text{odd on }1^{st})P(\text{odd on }2^{nd})\] \[=\frac{1}{2}\times\frac{1}{2}\]

OpenStudy (zarkon):

1,3,5 are odd numbers that are on the cube

OpenStudy (anonymous):

Thanks Zarkon. You got the answer, it was a simple calculation error. So it is (3/6)*(3/6) = 1/4